OpenStudy (anonymous):

the plane that contains the line x=-2+3t , y=4+2t , z=3-t and is perpandicular to the plane x-2y+z=5.

ganeshie8 (ganeshie8):

say, \(a(x-x_1) + b(y-y_1) + c(z-z_1)\) is the required plane

ganeshie8 (ganeshie8):

if this plane contains the given line, the point (-2, 4, 3) must satisfy the plane : \(a(x+2) + b(y-4)+c(z-3) = 0\)

ganeshie8 (ganeshie8):

Also, the direction vector of line and normal to plane must be perpendicular : \(3a + 2b - c = 0\) Since the plane is also perpendicular to another plane \(x-2y+z=5\), the normal vectors must be perpendicular : \(a - 2b + c = 0\) solving these two equations gives, \(<a, b, c> = <0, 1, 2>\)

ganeshie8 (ganeshie8):

that gives, \(0(x+2) + 1(y-4)+2(z-3) = 0\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

thanks alot :)

ganeshie8 (ganeshie8):

u wlc :)